OpenStudy (cas_fangirl_14):

Please help! Will medal. Algebra 1 B Question.

OpenStudy (cas_fangirl_14):

The probability of choosing a vowel (A, E, I, O, or U) from a deck of cards containing the 26 letters of the alphabet is shown below.  What is the probability of choosing the letters A and E one after the other without replacement? A. \[\frac{ 1 }{ 676 }\] B. \[\frac{ 1 }{ 650 }\] C. \[\frac{ 2 }{ 26 }\] D. \[\frac{ 1 }{ 52 }\]